Y has the following pdf. 20 2<p<4 ot herwise Find (a) 3 points F(1.8) (b) 3 points F(.5) (c) 3 points P(Y > 1.8 (d) 5 points E(Y) < 3.5) e) 5 points VO.

I have been able to get a,b,d, and e but can't find c.
a) F(1.8)=0.4358
b) F(3.5)=0.875
Answer #1

c) here P(Y>1.8|Y<3.5) =P(1.8<Y<3.5)/P(Y<3.5 )=(F(3.5)-F(1.8))/(F(3.5) =(0.875-0.4358)/0.875=0.519

( please revert for any clarification)